OpenStudy (anonymous):

Create a table for the exponential function f(x) = 3*2^x . My x values are ; -2 , -1, 0 , 2 , 1 .

OpenStudy (anonymous):

@tkhunny could u possibly help me with this one ?

OpenStudy (anonymous):

so find f(-2), f(-1), f(0), f(2), f(1)

OpenStudy (anonymous):

also, 1 should come before 2

OpenStudy (anonymous):

Lol sorry but yes that what it is . And the thing is idk how to find that

OpenStudy (anonymous):

@A_Harrold

OpenStudy (anonymous):

f = 3*2^x let x = -1 f= 3*(1/2) = 3/2 let x = 0 f = 3*2^0 = 3*1 = 3 etc.

OpenStudy (anonymous):

@Zale101